OpenStudy (anonymous):

Given that AE is a tangent of the circle with the center at B, BC = 3.5, and AE = 24, which is the length of EC?

OpenStudy (yanasidlinskiy):



OpenStudy (yanasidlinskiy):

You have a right triangle.... the chord AC is a diameter tangents and a radius intersect at right angles so angle EAB = 90 BC = 3.5 so AC = 7 (2 radii in a diameter) EC is found by Pythagoras' theorem \[EC^2=AE^2+AC^2\] \[EC^2=24^2+7^2\]